### v1

@@ 4.4 Linearization IV: Summary
$\bm{r}_{k}=-\textrm{discretization}\bigg{(}\frac{\partial\,\cdot\,}{\partial t}+\frac{\partial f(\,\cdot\,)}{\partial x}\bigg{)}\bar{\bm{u}}_{k},$ (55)
@@
$P_{k}\bar{\bm{e}}^{\textrm{lin}}_{k}=\textrm{discretization}\bigg{(}\frac{\partial\,\cdot\,}{\partial t}+\frac{\partial}{\partial x}\Big{(}f^{\prime}(\bar{\bm{u}}_{k})\,\cdot\,\Big{)}\bigg{)}\bar{\bm{e}}^{\textrm{lin}}_{k}=\bm{r}_{k}.$ (56)
@@
The linearization strategy outlined in this section is rather complicated, and a natural question is, does it need to be this complicated? For example, one could instead conceive of a simpler Picard iteration based on the quasi-linear form of the PDE (cons). Recall that wherever the solution of (cons) is differentiable, the PDE is equivalent to $\frac{\partial u}{\partial t}+f^{\prime}(u)\frac{\partial u}{\partial x}=0$. Given some approximation $u_{k}\approx u$, a residual correction scheme can then be formed by solving a linearized version of $\frac{\partial}{\partial t}(u_{k}+e_{k})+f^{\prime}(u_{k}+e_{k})\frac{\partial}{\partial x}(u_{k}+e_{k})=0$ arising from replacing the wave-speed with the approximation: $f^{\prime}(u_{k}+e_{k})\approx f^{\prime}(u_{k})$. At the discrete level, this leads to a linearized problem of the form
$\mathrm{discretization}\bigg{(}\frac{\partial\,\cdot\,}{\partial t}+f^{\prime}(\bar{\bm{u}}_{k})\frac{\partial\,\cdot\,}{\partial x}\bigg{)}\bar{\bm{e}}^{\mathrm{lin}}_{k}=\bm{r}_{k}.$ (57)
@@ 5 Approximate MGRIT solution of linearized problems
As detailed in Section 4, the blocks in the system matrix $P$ take the form of a certain, non-standard discretization of the linear conservation law (cons-lin). To simplify notation, in this section we write the linearized time-stepping operator $\Phi^{\textrm{lin}}(\bar{\bm{u}}^{n})$ as $\Phi^{n}$, where we recall from Section 3 that $\bar{\bm{u}}^{n}$ is the linearization point, and the linearized error $\bar{\bm{e}}^{\textrm{lin}}$ is written as $\bar{\bm{e}}$. [^4] Note, $\Phi^{n}$ is not to be confused with the nonlinear time-stepping operator $\Phi$ defined in Section 2, and the linearized error here $\bar{\bm{e}}$ is not to be confused with the true “nonlinear” error $\bar{\bm{e}}$ described in Section 3 even though they now use the same notation. Considering the structure of $P$ in (33), the $(n+1)$st block row of the system $P\bar{\bm{e}}=\bm{r}$ is $\bar{\bm{e}}^{n+1}-\Phi^{n}\bar{\bm{e}}^{n}=\bm{r}^{n+1}$.
@@ 5.1 MGRIT overview
Define a one-step coarse-grid matrix $P^{\Delta}\in\mathbb{R}^{\frac{n_{x}n_{t}}{m}\times\frac{n_{x}n_{t}}{m}}$ with the same block bidiagonal structure as $P$ in (33), such that the $(n+1)$st block row of its action on a coarse-grid vector $\bm{z}\in\mathbb{R}^{\frac{n_{x}n_{t}}{m}}$ is $\big{[}P^{\Delta}\bm{z}\big{]}^{n+1}=\bm{z}^{n+1}-\Psi^{n}\bm{z}^{n}$. Here, $\Psi^{n}$ is the (linear) coarse-grid time-stepping operator evolving vectors from time $t_{n}\to t_{n}+m\delta t$, and should be such that $\Psi^{n}\approx\Psi_{\rm ideal}^{n}:=\prod_{k=0}^{m-1}\Phi^{n+k}$, with $\Psi_{\rm ideal}^{n}$ being the ideal coarse-grid operator. Fast MGRIT convergence hinges on $\Psi^{n}$ being a sufficiently accurate approximation to $\Psi_{\rm ideal}^{n}$ [18, 59, 17], while obtaining speed-up in a parallel environment also necessitates that its action is significantly cheaper than that of $\Psi_{\rm ideal}^{n}$. As described in Section 1, the design of effective coarse-grid operators in the context of hyperbolic PDEs is a long-standing issue for MGRIT and Parareal.
@@ 5.2 Coarse-grid operator
$\Psi^{n}=\Big{(}I+{\cal T}_{\rm ideal}^{n}-{\cal T}_{\rm direct}^{n}\Big{)}^{-1}{\cal S}^{n,m\delta t}_{p,1_{*}}\approx\Psi_{\rm ideal}^{n}.$ (58)
Here, ${\cal S}^{n,m\delta t}_{p,1_{*}}$ is a conservative, FV-based discretization of (cons-lin) evolving solutions from time $t_{n}\to t_{n}+m\delta t$, and has, in a certain sense to be clarified below, orders of accuracy $p$ in space and one in time. We take $p$ as the formal order of the FV discretization described in Section 2, i.e., $p=2k-1$. Furthermore, ${\cal T}_{\rm direct}^{n}$ is an approximation of the operator associated with the largest term in the truncation error of ${\cal S}^{n,m\delta t}_{p,1_{*}}$. Further details on ${\cal S}^{n,m\delta t}_{p,1_{*}}$ and ${\cal T}_{\rm direct}^{n}$ are given in Section 5.3. Similarly, ${\cal T}_{\rm ideal}^{n}$ in (58) is an approximation of the operator associated with the largest truncation error term of $\Psi^{n}_{\rm ideal}$, and is discussed further below. The coarse-grid operator (58) has the same structure as the coarse-grid operators proposed in [14, 16] in the sense of being a semi-Lagrangian discretization modified so that its truncation error better approximates that of the ideal coarse-grid operator.
@@
Our approach for estimating ${\cal T}_{\rm ideal}^{n}$ is lengthy and details can be found in the supplementary materials. In essence, we first develop an approximate truncation error analysis in Supplementary Material Section SM4 for a linear ERK+FV-type discretization of the PDE (cons-lin). Then in Section SM5 this analysis is applied to estimate the truncation error of a standard linear ERK+FV discretization of (cons-lin). Numerical results are presented there to show that the approach results in a fast MGRIT solver despite the analysis being largely heuristic. This work generalizes [16], which considered (fine-grid) method-of-lines discretizations only for constant wave-speed problems. Finally, in Section SM6 the coarse-grid operator developed in Section SM5 is heuristically extended to the case at hand where the discretization of (cons-lin) corresponds to a linearization of an ERK+FV discretization of the nonlinear PDE (cons).
@@
${\cal T}_{\rm ideal}^{n}=\begin{cases}{\cal D}_{1}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\beta}^{n+j}\bigg{)}{\cal D}_{1}^{\top},\quad k=1,\\[12.91663pt] {\cal D}_{1}\bigg{(}\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\beta}^{1,n+j}\bigg{)}{\cal D}_{2}^{1}-\operatorname{diag}\bigg{(}\sum\limits_{j=0}^{m-1}\bm{\beta}^{0,n+j}\bigg{)}{\cal D}_{2}^{0}\bigg{)},\quad k=2.\end{cases}$ (59)
@@
$\beta_{i}^{n}$ $:=\left[-\frac{1}{2}\Big{(}\delta t\,\alpha^{n}_{i+1/2}\Big{)}^{2}+\frac{h\delta t}{2}\nu^{n}_{i+1/2}\right],$ (60a)
$\beta_{i}^{*,n}$ $:=\frac{3}{4}\left[s_{\rm RK}\,\frac{1}{24h}\Big{(}\delta t\,\alpha^{n}_{i+1/2}\Big{)}^{4}+s_{\rm FV}\,\frac{h^{2}\delta t}{12}\nu^{n}_{i+1/2}\right]\gamma_{i}^{*,n}.$ (60b)
In (5.2), $\alpha^{n}_{i+1/2}$ is an approximation to wave-speed $\alpha(x,t)$ of the linear conservation law (cons-lin) at $x=x_{i+1/2}$. Recalling from Section 4.4 that we are discretizing (cons-lin) with $\alpha(x,t)=f^{\prime}(u)$, we choose $\alpha^{n}_{i+1/2}=\frac{f^{\prime}(u_{i+1/2}^{-})+f^{\prime}(u_{i+1/2}^{+})}{2}$ in which $u_{i+1/2}^{\pm}$ are reconstructions of $u(x_{i+1/2})$ based on the either $\bar{\bm{u}}^{n}$ or $\bar{\bm{u}}^{n+1}$. Likewise, $\nu_{i+1/2}^{n}$ in (5.2) is the LFF dissipation coefficient in (6) arising from either $\bar{\bm{u}}^{n}$ or $\bar{\bm{u}}^{n+1}$.
@@
$\gamma_{i}^{0,n}:=2\widetilde{b}_{i+1}^{0,n}+b_{i}^{0,n},\quad\gamma_{i}^{1,n}:=\widetilde{b}_{i+1}^{1,n}+2b_{i}^{1,n},$ (61)
@@
In our numerical tests, we do not typically find a significant difference whether the time-dependent quantities in (5.2) are based on $\bar{\bm{u}}^{n}$ or $\bar{\bm{u}}^{n+1}$. An exception, however, is in our Buckley–Leverett test problems for order three (when $k=2$), where we find significantly better MGRIT convergence if the WENO weights, i.e., the $\{b_{i}^{\ell,n}\}$ and $\{\widetilde{b}_{i}^{\ell,n}\}$ in (61), are based on $\bar{\bm{u}}^{n+1}$ rather than $\bar{\bm{u}}^{n}$. As such, $k=2$ numerical tests in Section 6 (including for Burgers), compute all time-dependent quantities in (5.2) based on $\bar{\bm{u}}^{n+1}$; $k=1$ tests compute time-dependent quantities in (5.2) based on $\bar{\bm{u}}^{n}$.
@@
Finally, note that the truncation operator (59) is designed such that if optimal linear reconstruction weights are used in the discretization of (cons) are used rather than WENO weights then it reduces to the coarse-grid operator developed in Section SM5 for standard linear ERK+FV discretizations of (cons-lin).
@@ 5.3 Semi-Lagrangian discretization and its truncation error
Applying the discretization to (cons-lin) on the time interval $[t_{n},t_{n}+\delta t]$ results in the following approximate update for the cell average of $e(x)$ in cell ${\cal I}_{i}$:
$\bar{e}_{i}^{n+1}\approx\bar{e}_{i}^{n}-\frac{\widetilde{f}_{i+1/2}-\widetilde{f}_{i-1/2}}{h}=\big{(}{\cal S}_{p,r}^{n,\delta t}\bar{\bm{e}}^{n}\big{)}_{i}.$ (62)
@@
$\bar{e}_{i}^{n+1}=\frac{1}{h}\int_{\widetilde{\xi}_{i-1/2}}^{\widetilde{\xi}_{i+1/2}}e(x,t_{n})\operatorname{d}\!x=\bar{e}_{i}^{n}-\frac{1}{h}\left(\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}e(x,t_{n})\operatorname{d}\!x-\int_{\widetilde{\xi}_{i-1/2}}^{x_{i-1/2}}e(x,t_{n})\operatorname{d}\!x\right).$ (63)
@@
$\frac{\operatorname{d}\!\xi_{i+1/2}}{\operatorname{d}\!t}=\alpha(\xi_{i+1/2},t),\quad t\in[t_{n},t_{n}+\delta t],\quad\xi_{i+1/2}(t_{n}+\delta t)=x_{i+1/2}.$ (64)
@@
$\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}e(x,t_{n})\operatorname{d}\!x\approx\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}^{n}](x)\operatorname{d}\!x=:\widetilde{f}_{i+1/2}.$ (65)
Here, $R_{i+1/2}[\bar{\bm{e}}^{n}](x)\approx e(x,t_{n})$ is a cell-average-preserving piecewise polynomial reconstruction of $e(x,t_{n})$; see Section SM3.1 for further discussion.
@@
${\cal T}_{\rm direct}^{n}=-h^{p+1}{\cal D}_{1}\operatorname{diag}\big{(}g_{p+1}(\bm{\varepsilon})\big{)}{\cal D}_{p}^{\top}.$ (66)
This expression is based on a heuristic truncation error analysis for the semi-Lagrangian method applied to (cons-lin); further details can be found in Section SM3. Here, $\bm{\varepsilon}=\big{(}\varepsilon_{1+1/2},\ldots\varepsilon_{n_{x}+1/2}\big{)}\in\mathbb{R}^{n_{x}}$, with $\varepsilon_{i+1/2}$ the mesh-normalized difference between the $i$th departure point (over the coarse-grid time-step $t_{n}\to t_{n}+m\delta t$) and its east-neighboring mesh point; see the fine-grid example in Fig. 1. The function $g_{p+1}$ is a degree $p+1$ polynomial (see (SM18)) and is applied component-wise to $\bm{\varepsilon}$. As previously, $({\cal D}_{1}\bar{\bm{e}})_{i}=(\bar{e}_{i}-\bar{e}_{i-1})/h$, and ${\cal D}_{p}$, for $p$ odd, is a 1st-order accurate FD discretization of a $p$th derivative on a stencil with a 1-point bias to the left. Finally, we remark that (66) has a structure similar to the truncation error operator we used in [14], for FD discretizations of non-conservative, variable-wave-speed problems, which was of the form $-h^{p+1}\operatorname{diag}\big{(}g_{p+1}(\bm{\varepsilon})\big{)}{\cal D}_{p+1}$.

### v2

@@ 4.4 Linearization IV: Summary
(37) $\bm{r}_{k}=-\textrm{discretization}\bigg(\frac{\partial\,\cdot\,}{\partial t}+\frac{\partial f(\,\cdot\,)}{\partial x}\bigg)\bar{\bm{u}}_{k},$
@@
(38) $P_{k}\bar{\bm{e}}^{\textrm{lin}}_{k}=\textrm{discretization}\bigg(\frac{\partial\,\cdot\,}{\partial t}+\frac{\partial}{\partial x}\Big(f^{\prime}(\bar{\bm{u}}_{k})\,\cdot\,\Big)\bigg)\bar{\bm{e}}^{\textrm{lin}}_{k}=\bm{r}_{k}.$
@@
The linearization strategy outlined in this section is rather complicated, and it is natural to ask whether it needs to be this complicated. For example, one could instead conceive of a simpler Picard iteration based on the quasi-linear form of the PDE (cons). Recall that wherever the solution of (cons) is differentiable, the PDE is equivalent to $\frac{\partial u}{\partial t}+f^{\prime}(u)\frac{\partial u}{\partial x}=0$. Given some approximation $u_{k}\approx u$, a residual correction scheme can then be formed by solving a linearized version of $\frac{\partial}{\partial t}(u_{k}+e_{k})+f^{\prime}(u_{k}+e_{k})\frac{\partial}{\partial x}(u_{k}+e_{k})=0$ arising from replacing the wave-speed with the approximation: $f^{\prime}(u_{k}+e_{k})\approx f^{\prime}(u_{k})$. At the discrete level, this leads to a linearized problem of the form
(39) $\mathrm{discretization}\bigg(\frac{\partial\,\cdot\,}{\partial t}+f^{\prime}(\bar{\bm{u}}_{k})\frac{\partial\,\cdot\,}{\partial x}\bigg)\bar{\bm{e}}^{\mathrm{lin}}_{k}=\bm{r}_{k}.$
@@ 5 Approximate MGRIT solution of linearized problems
As detailed in Section 4, the blocks in the system matrix $P$ take the form of a certain, non-standard discretization of the linear conservation law (cons-lin). To simplify notation, in this section we write the linearized time-stepping operator $\Phi^{\textrm{lin}}(\bar{\bm{u}}^{n})$ as $\Phi^{n}$, where we recall from Section 3 that $\bar{\bm{u}}^{n}$ is the linearization point, and the linearized error $\bar{\bm{e}}^{\textrm{lin}}$ is written as $\bar{\bm{e}}$. [^4] Note, $\Phi^{n}$ is not to be confused with the nonlinear time-stepping operator $\Phi$ defined in Section 2, and the linearized error here $\bar{\bm{e}}$ is not to be confused with the true “nonlinear” error $\bar{\bm{e}}$ described in Section 3 even though they now use the same notation. Considering the structure of $P$ in (15), the (n+1)(n+1)st block row of the system $P\bar{\bm{e}}=\bm{r}$ is $\bar{\bm{e}}^{n+1}-\Phi^{n}\bar{\bm{e}}^{n}=\bm{r}^{n+1}$.
@@ 5.1 MGRIT overview
Define a coarse-grid space-time matrix $P^{\Delta}\in\mathbb{R}^{\frac{n_{x}n_{t}}{m}\times\frac{n_{x}n_{t}}{m}}$ with the same block bidiagonal structure as $P$ in (15), such that the (n+1)(n+1)st block row of its action on a coarse-grid vector $\bm{z}\in\mathbb{R}^{\frac{n_{x}n_{t}}{m}}$ is $\big[P^{\Delta}\bm{z}\big]^{n+1}=\bm{z}^{n+1}-\Psi^{n}\bm{z}^{n}$. Here, $\Psi^{n}$ is the (linear) coarse-grid time-stepping operator evolving vectors from time $t_{n}\to t_{n}+m\delta t$, and should be such that $\Psi^{n}\approx\Psi_{\rm ideal}^{n}:=\prod_{k=0}^{m-1}\Phi^{n+k}$, with $\Psi_{\rm ideal}^{n}$ being the ideal coarse-grid operator. Fast MGRIT convergence hinges on $\Psi^{n}$ being a sufficiently accurate approximation to $\Psi_{\rm ideal}^{n}$ [17, 59, 16], while obtaining speed-up in a parallel environment also necessitates that its action is significantly cheaper than that of $\Psi_{\rm ideal}^{n}$. As described in Section 1, the design of effective coarse-grid operators in the context of hyperbolic PDEs is a long-standing issue for MGRIT and Parareal.
@@ 5.2 Coarse-grid operator
(40) $\Psi^{n}=\Big(I+{\cal T}_{\rm ideal}^{n}-{\cal T}_{\rm direct}^{n}\Big)^{-1}{\cal S}^{n,m\delta t}_{p,1_{*}}\approx\Psi_{\rm ideal}^{n}.$
Here, ${\cal S}^{n,m\delta t}_{p,1_{*}}$ is a conservative, FV-based discretization of (cons-lin) evolving solutions from time $t_{n}\to t_{n}+m\delta t$, and has, in a certain sense to be clarified below, orders of accuracy $p$ in space and one in time. We take $p$ as the formal order of the FV discretization described in Section 2, i.e., p=2k1p=2k-1. Furthermore, ${\cal T}_{\rm direct}^{n}$ is an approximation of the operator associated with the leading-order term in the truncation error of ${\cal S}^{n,m\delta t}_{p,1_{*}}$. Further details on ${\cal S}^{n,m\delta t}_{p,1_{*}}$ and ${\cal T}_{\rm direct}^{n}$ are given in Section 5.3. Similarly, ${\cal T}_{\rm ideal}^{n}$ in (40) is an approximation of the operator associated with the leading-order term in the truncation error of $\Psi^{n}_{\rm ideal}$, and is discussed further below. The coarse-grid operator (40) has the same structure as the coarse-grid operators proposed in [14, 15] in the sense of being a semi-Lagrangian discretization modified so that its truncation error better approximates that of the ideal coarse-grid operator.
@@
Our approach for estimating ${\cal T}_{\rm ideal}^{n}$ is lengthy and details can be found in the supplementary materials. In essence, we first develop an approximate truncation error analysis in Supplementary Materials Section SM6 for a linear ERK+FV-type discretization of the PDE (cons-lin). Then in Section SM7 this analysis is applied to estimate the truncation error of a standard linear ERK+FV discretization of (cons-lin). Numerical results are presented there to show that the approach results in a fast MGRIT solver despite the analysis being largely heuristic. This work generalizes [15], which considered (fine-grid) method-of-lines discretizations only for constant wave-speed problems. Finally, in Section SM8 the coarse-grid operator developed in Section SM7 is heuristically extended to the case at hand where the discretization of (cons-lin) corresponds to a linearization of an ERK+FV discretization of the nonlinear PDE (cons).
@@
(41) ${\cal T}_{\rm ideal}^{n}=\begin{cases}{\cal D}_{1}\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\beta}^{n+j}\bigg){\cal D}_{1}^{\top},\quad k=1,\\[12.91663pt] {\cal D}_{1}\bigg(\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\beta}^{1,n+j}\bigg){\cal D}_{2}^{1}-\operatorname{diag}\bigg(\sum\limits_{j=0}^{m-1}\bm{\beta}^{0,n+j}\bigg){\cal D}_{2}^{0}\bigg),\quad k=2.\end{cases}$
@@
(42a) $\beta_{i}^{n}$ $:=\left[-\frac{1}{2}\Big(\delta t\,\alpha^{n}_{i+1/2}\Big)^{2}+\frac{h\delta t}{2}\nu^{n}_{i+1/2}\right],$
(42b) $\beta_{i}^{*,n}$ $:=\frac{3}{4}\left[s_{\rm RK}\,\frac{1}{24h}\Big(\delta t\,\alpha^{n}_{i+1/2}\Big)^{4}+s_{\rm FV}\,\frac{h^{2}\delta t}{12}\nu^{n}_{i+1/2}\right]\gamma_{i}^{*,n}.$
In (42), $\alpha^{n}_{i+1/2}$ is an approximation to wave-speed $\alpha(x,t)$ of the linear conservation law (cons-lin) at $x=x_{i+1/2}$. Recalling from Section 4.4 that we are discretizing (cons-lin) with $\alpha(x,t)=f^{\prime}(u)$, we choose $\alpha^{n}_{i+1/2}=\frac{f^{\prime}(u_{i+1/2}^{-})+f^{\prime}(u_{i+1/2}^{+})}{2}$ in which $u_{i+1/2}^{\pm}$ are reconstructions of $u(x_{i+1/2})$ based on the either $\bar{\bm{u}}^{n}$ or $\bar{\bm{u}}^{n+1}$. Likewise, $\nu_{i+1/2}^{n}$ in (42) is the LFF dissipation coefficient in (6) arising from either $\bar{\bm{u}}^{n}$ or $\bar{\bm{u}}^{n+1}$.
@@
(43) $\gamma_{i}^{0,n}:=2\widetilde{b}_{i+1}^{0,n}+b_{i}^{0,n},\quad\gamma_{i}^{1,n}:=\widetilde{b}_{i+1}^{1,n}+2b_{i}^{1,n},$
@@
In our numerical tests, we do not typically find a significant difference whether the time-dependent quantities in (42) are based on $\bar{\bm{u}}^{n}$ or $\bar{\bm{u}}^{n+1}$. An exception, however, is in our Buckley–Leverett test problems for order three (when $k=2$), where we find significantly better MGRIT convergence if the WENO weights, i.e., the $\{b_{i}^{\ell,n}\}$ and $\{\widetilde{b}_{i}^{\ell,n}\}$ in (43), are based on $\bar{\bm{u}}^{n+1}$ rather than $\bar{\bm{u}}^{n}$. As such, $k=2$ numerical tests in Section 6 (including for Burgers), compute all time-dependent quantities in (42) based on $\bar{\bm{u}}^{n+1}$; $k=1$ tests compute time-dependent quantities in (42) based on $\bar{\bm{u}}^{n}$.
@@
Finally, note that the truncation operator (41) is designed such that if optimal linear reconstruction weights are used in the discretization of (cons) are used rather than WENO weights then it reduces to the coarse-grid operator developed in Supplementary Materials Section SM7 for standard linear ERK+FV discretizations of (cons-lin).
@@ 5.3 Semi-Lagrangian discretization and its truncation error
Applying the discretization to (cons-lin) on the time interval $t_{n},t_{n}+\delta t]$ results in the following approximate update for the cell average of $e(x)$ in cell ${\cal I}_{i}$:
(44) $\bar{e}_{i}^{n+1}\approx\bar{e}_{i}^{n}-\frac{\widetilde{f}_{i+1/2}-\widetilde{f}_{i-1/2}}{h}=\big({\cal S}_{p,r}^{n,\delta t}\bar{\bm{e}}^{n}\big)_{i}.$
@@
(45) $\bar{e}_{i}^{n+1}=\frac{1}{h}\int_{\widetilde{\xi}_{i-1/2}}^{\widetilde{\xi}_{i+1/2}}e(x,t_{n})\operatorname{d}\!x=\bar{e}_{i}^{n}-\frac{1}{h}\left(\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}e(x,t_{n})\operatorname{d}\!x-\int_{\widetilde{\xi}_{i-1/2}}^{x_{i-1/2}}e(x,t_{n})\operatorname{d}\!x\right).$
@@
(46) $\frac{\operatorname{d}\!\xi_{i+1/2}}{\operatorname{d}\!t}=\alpha(\xi_{i+1/2},t),\quad t\in[t_{n},t_{n}+\delta t],\quad\xi_{i+1/2}(t_{n}+\delta t)=x_{i+1/2}.$
@@
(47) $\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}e(x,t_{n})\operatorname{d}\!x\approx\int_{\widetilde{\xi}_{i+1/2}}^{x_{i+1/2}}R_{i+1/2}[\bar{\bm{e}}^{n}](x)\operatorname{d}\!x=:\widetilde{f}_{i+1/2}.$
Here, $R_{i+1/2}[\bar{\bm{e}}^{n}](x)\approx e(x,t_{n})$ is a cell-average-preserving piecewise polynomial reconstruction of $e(x,t_{n})$; see Supplementary Materials Section SM5.1 for further discussion.
@@
(48) ${\cal T}_{\rm direct}^{n}=-h^{p+1}{\cal D}_{1}\operatorname{diag}\big(g_{p+1}(\bm{\varepsilon})\big){\cal D}_{p}^{\top}.$
This expression is based on a heuristic truncation error analysis for the semi-Lagrangian method applied to (cons-lin); further details can be found in Supplementary Materials Section SM5. Here, $\bm{\varepsilon}=\big(\varepsilon_{1+1/2},\ldots\varepsilon_{n_{x}+1/2}\big)\in\mathbb{R}^{n_{x}}$, with $\varepsilon_{i+1/2}$ the mesh-normalized difference between the $i$th departure point (over the coarse-grid time-step $t_{n}\to t_{n}+m\delta t$) and its east-neighboring mesh point; see the fine-grid example in Fig. 1. The function $g_{p+1}$ is a degree $p+1$ polynomial (see (SM30)) and is applied component-wise to $\bm{\varepsilon}$. As previously, $({\cal D}_{1}\bar{\bm{e}})_{i}=(\bar{e}_{i}-\bar{e}_{i-1})/h$, and ${\cal D}_{p}$, for $p$ odd, is a 1st-order accurate FD discretization of a $p$th derivative on a stencil with a 1-point bias to the left. Finally, we remark that (48) has a structure similar to the truncation error operator we used in [14], for FD discretizations of non-conservative, variable-wave-speed problems, which was of the form $h^{p+1}\operatorname{diag}\big(g_{p+1}(\bm{\varepsilon})\big){\cal D}_{p+1}$.